# Targeted Nakamoto: A Bitcoin Protocol to Balance Network Security and Energy Consumption

### v2

@@ Abstract
footnotetext: I wish to thank Madars Virza for his aid and advice in developing the core ideas of this paper; Neha Narula for helping to refine and clarify the presentation; thoughtful comments from participants at the Chaincode Labs research seminar where an early version of this paper was presented, and Mark Serrahn for producing the figures herein.
@@ 1 Introduction
Targeted Nakamoto (hereafter ”Targeted”) modifies Nakamoto to create an incentive for miners to push hashrate toward an interval whenever it strays outside the interval. The miner’s minimum block reward is increased when hashrate is below the interval, which induces miners to apply more hashrate. The miner’s maximum block reward is decreased when hashrate is above the interval, which induces miners to apply less hashrate. Monetary neutrality is maintained by offsetting the increase and decrease in the miner’s block reward by opposite adjustments to the spending potential of addresses with UTXO value, in proportion to the value held by each address. Targeted mitigates two latent, potentially existential threats to Bitcoin. One threat arises from the halving of the minting of UTXO’s (the ”coinbase”), which ensures that transaction fees will eventually become the primary source of block rewards. [^5] Historically, the trend USD/BTC exchange rate has risen sufficiently to offset the halving of the coinbase. However, this cannot continue indefinitely for two reasons. First, coinbase will eventually be zero. Second, maintaining a fixed USD value of the coinbase paid to miners as halving continues would require the inflation adjusted USD to BTC exchange rate to double every four years. This implies a growth rate that exceeds US (and world) GDP by orders of magnitude, which is dynamically unsustainable (Tirole [Tir85]). A transactor is motivated to add a fee to incentivize a miner to include the transaction in its block. [^6] Huberman et.al. [HLM21] A transactor has, at best, a muted incentive to include a fee for network security, which is a non-rival public good that benefits all participants. Historically, transaction fees have comprised less than 10% of total block rewards on average and there is no assurance that transaction fees will increase in the future. [^7] Data on USD block rewards and transaction fees for Bitcoin are compiled by Coinmetrics https://coinmetrics.io The scheduled decline in the coinbase, if not matched by an increase in transaction fees, will cause miners to reduce hashrate, which will lower the cost of an attack on the network. The other threat comes from the reality and the adverse political reaction to the fact that the electricity consumed by Bitcoin miners adds to carbon emissions. Targeted mitigates both of these risks by incentivizing miners to home in on a hashrate interval. This ensures that the cost of attacking the network will not become dangerously cheap while limiting the network’s contribution to carbon emissions and electricity cost. Finally, Targeted preserves aggregate monetary neutrality by proportionally adjusting the spending potential of Bitcoin addresses to offset the minting or burning of UTXO value that that arises from the provision of incentives to miners.
@@ 1.1 Limitation
The decrease in network security risk and the increase in electricity cost and carbon emissions as hashrate increases ensures, with a mild restriction on curvature, the existence of a point or interval at which the sum of the direct and externalities costs are minimized. [^8] The restriction is that the function defined as the sum of the costs is is concave. However, locating the minimum requires a numerical assignment of the costs of network security risk and carbon emissions, which lies outside the scope of this paper. A related issue is the tracking of changes in the relationship between hashrate and electricity cost. The efficiency of Bitcoin ASICS has improved over time, reducing the Kw per hash. Coinmetrics [Coi23] estimates that the average electricity per hash among active Bitcoin ASIC’s decreased by half from 2018 to 2022. If improvements continue, a fixed target hashrate interval will be associated with lower electricity consumption and a lower cost of attack. The difficulty here is not the tracking of changes in per kw electricity cost and ASIC efficiency, but rather use of an oracle that it requires, which violates a value held by many in the Bitcoin community. Finally, the relation between electricity cost and security may be non-stationary. A change in the USD value of Bitcoin, or in the size distribution or other characteristics of transactions, may alter the security achieved at any level of hashrate. Against these limitations stand the existential threats mentioned above, which may convince the reader of the need to form consensus on a target interval notwithstanding the gaps in our knowledge.
@@ 1.2 Related Literature
This paper builds on the contributions of Nakamoto [Nak08] who laid the foundations for the PoW protocol used in Bitcoin.
@@
There are two previous works with which aspects of this study closely align. Pagnotta (2021) [Pag21] identified and modeled the complimentarity between mining hashrate and Bitcoin USD price. The interdependence runs through network security. Hashrate determines security, security affects traders valuation of Bitcoin, which affects it price, price affects the USD value of the mining reward, which determines equilibrium hashrate. Targeted injects into this causal chain a regulator that adjusts the miner’s block reward to incentivize miners to keep hashrate (and thereby security) in the neighborhood of a selected target. Mirkin et.al. (2024) [MZEZ24] show that electricity cost can be reduced, for any level of network security, by adding a ”proof-of-delay” (e.g. using a verifiable delay function) threshold for a miner’s block to be eligible to append to the chain. The mandatory delay interval creates an incentive for miners to shift expenditure from hashrate to investment in mining equipment. The re-allocation of expenditure from variable to fixed cost reduces the carbon footprint. Notably, proof-of-delay does not alter equilibrium expenditure on mining, which is determined by the miner’s block reward. This means it does not affect the cost to attack the network and thereby network security. While proof-of-delay reduces energy consumption for a given level of security, Targeted, which is compatible with proof-of-delay, enables network security to be selected and maintained.
@@
One of the externalities costs addressed in this paper are the carbon emissions generated by the electricity consumed in mining Bitcoin. I rely on two data providers. One is the Bitcoin Electricity Consumption Index (BECI) of the Cambridge U.K. Centre for Alternative Finance [fAF24], which is regarded as a credible - though not infallible - source for information on Bitcoin energy consumption. The other is Coinmeterics  [^9] https://coinmetrics.io(which is a source for BECI) for other Bitcoin performance indicators. Coinmetrics is considered to be a careful aggregator of data on cryptocurrencies. Sedlmeir et.al. [SUFK20] and Gallersdorfer et.al. [GKS20] review comparative magnitudes of energy consumption for different blockchain protocols. Gallersdorfer [GKS20] compares the estimation of Bitcoin electricity consumption across several sources. For my purposes, the salient finding is that these studies document a time-path of energy consumption that is tightly correlated with the time-path of hashrate in Bitcoin.  [^10] See Figure 1 in Gellersdorfer [GKS20].
@@
As with several previous studies of blockchain protocol, this paper uses mechanism design approach to the study of the consensus protocol. Chen et.al. [LWX20] surveys various aspects of existing PoW blockchain protocols, including the incentive to increase energy consumption in mining, from a mechanism design perspective. Cong et.al [CHL21] analyses the effect of PoW mining competition on energy consumption. Leshno and Strack [LS20] identifies features of the miner payoff structure that are required to maintain blockchain stability. They prove that certain of the core properties of a decentralized PoW blockchain - namely free entry of anonymous profit driven miners - are only incentive compatible with a reward mechanism that is similar to Nakamoto. Targeted preserves those features.
@@
In their study of the Bitcoin payment system, Huberman et.al. [HLM21] characterize an equilibrium where transactions are appended to blocks in descending order of the size of fee offered by transaction senders, which reflects transactor’s relative willingness to pay for a position in the queue for placement on the blockchain. The adjustment to the miner’s block reward in Targeted is guided by this insight. When the block reward is adjusted, the ordering of the rewards received by the miner from transactions always matches the order of fees offered by transaction senders. This, in turn, leaves unaltered he order in which transactions are appended to blocks.
@@ 1.3 The research question and design constraints
It will be shown that Targeted satisfies each of these constraints with one exception. Targeted can require an increase in the coinbase, which is a hard-fork.
@@ 2.1 Externalities costs that increase with hashrate
Hashing, or puzzle guessing, involves computation that consumes electricity. This generates two external effects. One is to increase the price of electricity and the other is to increase the social cost of carbon emissions from the from fossil fuel burning electricity generators. [^12] Chen et.al. [LWX20] model the flow of carbon emissions from multiple energy sources. These costs are related in the sense that an increase in hashrate will, by increasing electricity demand, put upward pressure on electricity price and, to the extent that the increase in demand is met by an increase in electricity generation (i.e. that electricity supply is not inelastic), cause a rise in carbon emissions. [^13] Fowlie et.al. [FRR16] estimate a price elasticity of demand for electricity in the U.S. of -2.06 for with a range of alternative elasticities around -1 (see Table 2). These results show that an increase in electricity demand will be split between increases in both price and quantity.
@@ 2.2 Externalities costs that decrease with hashrate
Double-spend attack A double-spend attack requires the attacker to build an alternative chain that overtakes the incumbent canonical chain at in terms of cumulative mining puzzle difficulty. An attacker with over 50% of total hashrate will overtake the incumbent canonical chain at some block with probability measure 1. At less than 50% share, the attacker’s probability of success is less than half. [^14] See Gervais et.al. [GGK+16] for simulations of attacks using different ratios of attacker hashrate to miner hashrate. The frequency of success is increasing in attacker hashrate and the average number of blocks required to overtake the incumbent canonical chain is decreasing in attacker hashrate. The amount of attacker hashrate required to successfully fork the blockchain, and its cost, is increasing in the level of honest miner hashrate. Budish (2018) [Bud18] and Aronoff et.al. [AA24] point out that the cost of a successful attack is at least partially offset by the block rewards that accrue to the attacker. The possibility that a victim of a double-spend attack retaliates by mining on the incumbent canonical chain leads to a war of attrition between the attacker and the victim (Moroz et.al. [MANP20]). Aronoff et.al. [AA24]) point out that the outcome of a war of attrition depends on equilibrium refinements and exigent circumstances, even when the costs between players differ. Nevertheless, an increase in the cost of launching the attack will act as a deterrent by increasing the sunk cost that the attacker loses if it is unsuccessful. Equation 1 is the attacker’s un-discounted ex-ante expected profit when block rewards equal hashrate cost. The key observation is that, for any probability of success, the attacker’s expected profit is declining in honest miner hashrate.
@@ 2.3 The target hashrate interval
Figure 2 is a conceptual depiction of the curves representing network security and energy externalities cost. The horizontal axis is the hashrate of miners in the Bitcoin network. The vertical axis represents the cost of security and energy externalities. Total cost is the summation of these curves at each level of hashrate. The goal of Targeted is minimize the total cost of security and energy externalities. The negative slope of the network security curve reflects that the network becomes less vulnerable as hashrate increases. Network security vanishes as hashrate approaches zero, since the cost of acquiring 51% hashrate is negligible (equivalent to unbounded security cost). The cost of attacking the network becomes extremely expensive at very high levels (equivalent to negligible security cost). A mild smoothness assumption implies that the network security curve is convex in hashrate. I assume the unit hashrate cost is either constant or increases with electricity consumption (and hashrate). The shape of the social cost of carbon emissions is convex in electricity consumption (and hashrate).161616 ”The best evidence is that impacts of climate change will be non-linear and cumulative.” (William Nordhaus Nobel Prize lecture in economics [Nor19]). Deitz et.al. [DSN15] identify the ”three essential elements of the climate problem – the endogeneity of growth, the convexity of damage and climate risk”. The energy externality cost is the summation of these two costs, which is convex. The sum of the two curves is convex, which assures the existence of an interval in which total cost is minimized. A reduction in per hash enery cost induced by improvements in ASIC efficiency and/or electricity cost will shift the network security curve to the right. A reduction in per hash carbon emissions from an increase in the use of clean energy will shift the hashrate/energy externalities cost curve to the right.
@@ 3.2 An overview of the mechanism design of the protocols
Target Blockchain Growth Rate Hashrate Upper Bound
@@ 4.1 The puzzle difficulty signal of hashrate
A single mining computer can guess a puzzle solution sequentially in time. [^19] A computer with $n$ parallel processors is treated here as $n$ separate computers. I model mining computers as making successive guesses synchronously at times $t\in\{t_{1},t_{2},...\}$, each separated by an interval of temporal length $t_{i+1}-t_{i}$ which I normalize to 1 unit of time [^20] The synchronicity enables the expression of the time-path of puzzle solutions compactly as a Bernoulli Process, which simplifies the analysis WLOG.. The Nakamoto blockchain growth target is one block every $T^{*}$ units of time. During this time interval each computer has guessed (or hashed) $T^{*}$ proposed solutions. I define a unit of hashrate as a single mining computer making $T^{*}$ puzzle guesses. If there are $N$ mining computers making guesses, the number of total guesses made in the time interval of the blockchain growth target is $NT^{*}$ and the hashrate is $N$212121 This accords with the conventional definition of hashrate used e.g. by Coinmetrics [Coi20], as puzzle guesses per second. The framework for guessing puzzle solutions over time is a Bernoulli Process. Let $T$ denote the time the first puzzle solution is broadcast to the network. For a hashrate of $N$ the expected time to reach the solution is expressed by Equation 2.

### v3

@@ Abstract
Bitcoin faces two, diametrically opposed and temporally divergent, existential risks. Today, the massive energy consumed by Bitcoin mining is a focus of a political backlash aimed at de-legitimizing Bitcoin and restricting its use. In the future, as the minting of new UTXO’s to reward miners decreases toward zero, the revenue from mining will likely shrink, causing miners to exit and energy consumption to contract. The security of the network may be compromised by the low cost of an attack. In effect, Bitcoin’s energy consumption will be too low. Targeted Nakamoto is a modification to the Bitcoin protocol that is designed to avert these twin potential disasters.
footnotetext: I wish to thank Madars Virza for his aid and advice in developing the core ideas of this paper; Neha Narula for helping to refine and clarify the presentation; thoughtful comments from participants at the Chaincode Labs research seminar and Mark Serrahn for producing the figures herein. Kristain Praizner wrote the code for the hashrate control algorithm and implemented it on an API that is a companion to this paper.
@@ 1 Introduction
Targeted Nakamoto (hereafter ”Targeted”) modifies Nakamoto to create an incentive for miners to push hashrate toward an interval whenever it strays outside the interval. The miner’s minimum block reward is increased when hashrate is below the interval, which induces miners to apply more hashrate. The miner’s maximum block reward is decreased when hashrate is above the interval, which induces miners to apply less hashrate. Monetary neutrality is maintained by offsetting the increase and decrease in the miner’s block reward by opposite adjustments to the spending potential of addresses with UTXO value, in proportion to the value held by each address.
1.1 Two existential risks to Bitcoin
@@
Targeted mitigates two latent, potentially existential threats to Bitcoin. One threat arises from the halving of the minting of UTXO’s (the ”coinbase”), which ensures that transaction fees will eventually become the primary source of block rewards. [^5] Historically, the trend USD/BTC exchange rate has risen sufficiently to offset the halving of the coinbase. However, this cannot continue indefinitely for two reasons. First, coinbase will eventually be zero. Second, maintaining a fixed USD value of the coinbase paid to miners as halving continues would require the inflation adjusted USD to BTC exchange rate to double every four years. This implies a growth rate that exceeds US (and world) GDP by orders of magnitude, which is dynamically unsustainable (Tirole [21]). A transactor is motivated to add a fee to incentivize a miner to include the transaction in its block. [^6] Huberman et.al. [12] A transactor has, at best, a muted incentive to include a fee for network security, which is a non-rival public good that benefits all participants. Historically, transaction fees have comprised less than 10% of total block rewards on average and there is no assurance that transaction fees will increase in the future. [^7] Data on USD block rewards and transaction fees for Bitcoin are compiled by Coinmetrics https://coinmetrics.io The scheduled decline in the coinbase, if not matched by an increase in transaction fees, will cause miners to reduce hashrate, which will lower the cost of an attack on the network. The other threat comes from the reality and the adverse political reaction to the fact that the electricity consumed by Bitcoin miners adds to carbon emissions. Targeted mitigates both of these risks by incentivizing miners to home in on a hashrate interval. This ensures that the cost of attacking the network will not become dangerously cheap while limiting the network’s contribution to carbon emissions and electricity cost. Finally, Targeted preserves aggregate monetary neutrality by proportionally adjusting the spending potential of Bitcoin addresses to offset the minting or burning of UTXO value that that arises from the provision of incentives to miners.
1.2 Links to code and web API
Accompanying this paper is a link to a code-base for controlling Hashrate and a Web API with an empirically estimated hashrate function and a module that enables the user to choose a target hashrate interval and control parameters and to run counterfactual experiments with alternative time-paths of model variables. Both instantiations were developed by Kristian Praizner of MIT.
The code-base is here https://github.com/Krisp140/TargetedNakomoto
The Web API is here https://targetednakamoto.com/About.
Both are anonymized for this submission.
@@ 1.3 Limitation
The decrease in network security risk and the increase in electricity cost and carbon emissions as hashrate increases ensures, with a mild restriction on curvature, the existence of a point or interval at which the sum of the direct and externalities costs are minimized. [^8] The restriction is that the function defined as the sum of the costs is is concave. However, locating the minimum requires a numerical assignment of the costs of network security risk and carbon emissions, which lies outside the scope of this paper. A related issue is the tracking of changes in the relationship between hashrate and electricity cost. The efficiency of Bitcoin ASICS has improved over time, reducing the Kw per hash. Coinmetrics [5] estimates that the average electricity per hash among active Bitcoin ASIC’s decreased by half from 2018 to 2022. If improvements continue, a fixed target hashrate interval will be associated with lower electricity consumption and a lower cost of attack. The difficulty here is not the tracking of changes in per kw electricity cost and ASIC efficiency, but rather use of an oracle that it requires, which violates a value held by many in the Bitcoin community. Finally, the relation between electricity cost and security may be non-stationary. A change in the USD value of Bitcoin, or in the size distribution or other characteristics of transactions, may alter the security achieved at any level of hashrate. Against these limitations stand the existential threats mentioned above, which may convince the reader of the need to form consensus on a target interval notwithstanding the gaps in our knowledge.
@@ 1.4 Related Literature
This paper builds on the contributions of Nakamoto [17] who laid the foundations for the PoW protocol used in Bitcoin.
@@
There are two previous works with which aspects of this study closely align. Pagnotta [19] identified and modeled the complimentarity between mining hashrate and Bitcoin USD price. The interdependence runs through network security. Hashrate determines security, security affects traders valuation of Bitcoin, which affects it price, price affects the USD value of the mining reward, which determines equilibrium hashrate. Targeted injects into this causal chain a regulator that adjusts the miner’s block reward to incentivize miners to keep hashrate (and thereby security) in the neighborhood of a selected target. Mirkin et.al. [15] show that electricity cost can be reduced, for any level of network security, by adding a ”proof-of-delay” (e.g. using a verifiable delay function) threshold for a miner’s block to be eligible to append to the chain. The mandatory delay interval creates an incentive for miners to shift expenditure from hashrate to investment in mining equipment. The re-allocation of expenditure from variable to fixed cost reduces the carbon footprint. Notably, proof-of-delay does not alter equilibrium expenditure on mining, which is determined by the miner’s block reward. This means it does not affect the cost to attack the network and thereby network security. While proof-of-delay reduces energy consumption for a given level of security, Targeted, which is compatible with proof-of-delay, enables network security to be selected and maintained.
@@
One of the externalities costs addressed in this paper are the carbon emissions generated by the electricity consumed in mining Bitcoin. I rely on two data providers. One is the Bitcoin Electricity Consumption Index (BECI) of the Cambridge U.K. Centre for Alternative Finance [8], which is regarded as a credible - though not infallible - source for information on Bitcoin energy consumption. The other is Coinmeterics  [^9] https://coinmetrics.io(which is a source for BECI) for other Bitcoin performance indicators. Coinmetrics is considered to be a careful aggregator of data on cryptocurrencies. Sedlmeir et.al. [20] and Gallersdorfer et.al. [10] review comparative magnitudes of energy consumption for different blockchain protocols. Gallersdorfer [10] compares the estimation of Bitcoin electricity consumption across several sources. For my purposes, the salient finding is that these studies document a time-path of energy consumption that is tightly correlated with the time-path of hashrate in Bitcoin.  [^10] See Figure 1 in Gellersdorfer [10].
@@
As with several previous studies of blockchain protocol, this paper uses mechanism design approach to the study of the consensus protocol. Chen et.al. [14] surveys various aspects of existing PoW blockchain protocols, including the incentive to increase energy consumption in mining, from a mechanism design perspective. Cong et.al [6] analyses the effect of PoW mining competition on energy consumption. Leshno and Strack [13] identifies features of the miner payoff structure that are required to maintain blockchain stability. They prove that certain of the core properties of a decentralized PoW blockchain - namely free entry of anonymous profit driven miners - are only incentive compatible with a reward mechanism that is similar to Nakamoto. Targeted preserves those features.
@@
In their study of the Bitcoin payment system, Huberman et.al. [12] characterize an equilibrium where transactions are appended to blocks in descending order of the size of fee offered by transaction senders, which reflects transactor’s relative willingness to pay for a position in the queue for placement on the blockchain. The adjustment to the miner’s block reward in Targeted is guided by this insight. When the block reward is adjusted, the ordering of the rewards received by the miner from transactions always matches the order of fees offered by transaction senders. This, in turn, leaves unaltered he order in which transactions are appended to blocks.
@@ 1.5 The research question and design constraints
It will be shown that Targeted satisfies each of these constraints with one exception. Targeted can instruct miners to burn UTXO’s from transaction outputs, which may require a hard-fork to implement.
@@ 2.1 Externalities costs that increase with hashrate
Hashing, or puzzle guessing, involves computation that consumes electricity. This generates two external effects. One is to increase the price of electricity and the other is to increase the social cost of carbon emissions from the from fossil fuel burning electricity generators. [^12] Chen et.al. [14] model the flow of carbon emissions from multiple energy sources. These costs are related in the sense that an increase in hashrate will, by increasing electricity demand, put upward pressure on electricity price and, to the extent that the increase in demand is met by an increase in electricity generation (i.e. that electricity supply is not inelastic), cause a rise in carbon emissions. [^13] Fowlie et.al. [9] estimate a price elasticity of demand for electricity in the U.S. of -2.06 for with a range of alternative elasticities around -1 (see Table 2). These results show that an increase in electricity demand will be split between increases in both price and quantity.
@@ 2.2 Externalities costs that decrease with hashrate
Double-spend attack A double-spend attack requires the attacker to build an alternative chain that overtakes the incumbent canonical chain at in terms of cumulative mining puzzle difficulty. An attacker with over 50% of total hashrate will overtake the incumbent canonical chain at some block with probability measure 1. At less than 50% share, the attacker’s probability of success is less than half. [^14] See Gervais et.al. [11] for simulations of attacks using different ratios of attacker hashrate to miner hashrate. The frequency of success is increasing in attacker hashrate and the average number of blocks required to overtake the incumbent canonical chain is decreasing in attacker hashrate. The amount of attacker hashrate required to successfully fork the blockchain, and its cost, is increasing in the level of honest miner hashrate. Budish (2018) [3] and Aronoff et.al. [1] point out that the cost of a successful attack is at least partially offset by the block rewards that accrue to the attacker. The possibility that a victim of a double-spend attack retaliates by mining on the incumbent canonical chain leads to a war of attrition between the attacker and the victim (Moroz et.al. [16]). Aronoff et.al. [1]) point out that the outcome of a war of attrition depends on equilibrium refinements and exigent circumstances, even when the costs between players differ. Nevertheless, an increase in the cost of launching the attack will act as a deterrent by increasing the sunk cost that the attacker loses if it is unsuccessful. Equation 1 is the attacker’s un-discounted ex-ante expected profit when block rewards equal hashrate cost. The key observation is that, for any probability of success, the attacker’s expected profit is declining in honest miner hashrate.
@@ 2.3 The target hashrate interval
Figure 2 is a conceptual depiction of the curves representing network security and energy externalities cost. The horizontal axis is the hashrate of miners in the Bitcoin network. The vertical axis represents the cost of security and energy externalities. Total cost is the summation of these curves at each level of hashrate. The goal of Targeted is minimize the total cost of security and energy externalities. The negative slope of the network security curve reflects that the network becomes less vulnerable as hashrate increases. Network security vanishes as hashrate approaches zero, since the cost of acquiring 51% hashrate is negligible (equivalent to unbounded security cost). The cost of attacking the network becomes extremely expensive at very high levels (equivalent to negligible security cost). A mild smoothness assumption implies that the network security curve is convex in hashrate. I assume the unit hashrate cost is either constant or increases with electricity consumption (and hashrate). The shape of the social cost of carbon emissions is convex in electricity consumption (and hashrate).161616 ”The best evidence is that impacts of climate change will be non-linear and cumulative.” (William Nordhaus Nobel Prize lecture in economics [18]). Deitz et.al. [7] identify the ”three essential elements of the climate problem – the endogeneity of growth, the convexity of damage and climate risk”. The energy externality cost is the summation of these two costs, which is convex. The sum of the two curves is convex, which assures the existence of an interval in which total cost is minimized. A reduction in per hash enery cost induced by improvements in ASIC efficiency and/or electricity cost will shift the network security curve to the right. A reduction in per hash carbon emissions from an increase in the use of clean energy will shift the hashrate/energy externalities cost curve to the right.
@@ 3.2 An overview of the mechanism design of the protocols
Target Blockchain Growth Rate Hashrate Target Interval
@@ 4.1 The puzzle difficulty signal of hashrate
A single mining computer can guess a puzzle solution sequentially in time. [^19] A computer with $n$ parallel processors is treated here as $n$ separate computers. I model mining computers as making successive guesses synchronously at times $t\in\{t_{1},t_{2},...\}$, each separated by an interval of temporal length $t_{i+1}-t_{i}$ which I normalize to 1 unit of time [^20] The synchronicity enables the expression of the time-path of puzzle solutions compactly as a Bernoulli Process, which simplifies the analysis WLOG.. The Nakamoto blockchain growth target is one block every $T^{*}$ units of time. During this time interval each computer has guessed (or hashed) $T^{*}$ proposed solutions. I define a unit of hashrate as a single mining computer making $T^{*}$ puzzle guesses. If there are $N$ mining computers making guesses, the number of total guesses made in the time interval of the blockchain growth target is $NT^{*}$ and the hashrate is $N$212121 This accords with the conventional definition of hashrate used e.g. by Coinmetrics [4], as puzzle guesses per second. The framework for guessing puzzle solutions over time is a Bernoulli Process. Let $T$ denote the time the first puzzle solution is broadcast to the network. For a hashrate of $N$ the expected time to reach the solution is expressed by Equation 2.